OpenStudy (anonymous):

Luke filled the glasses shown below completely with water. The total amount of water that Luke poured into the glasses is 90 cubic centimeters. What is the height of glass 1? Round your answer to the nearest hundredths. (Use π = 3.14.) Note that all measurements are in centimeters and images are not drawn to scale. A cylinder with width 4.5 and height unknown is labeled glass 1, and a cone with height 5.6 and width 4.2 is labeled glass 2. 2.63 centimeters 3.82 centimeters 4.04 centimeters 5.62 centimeters

OpenStudy (anonymous):

@mathmate @Elsa213

OpenStudy (mathmate):

Are the glasses filled to the same height (h) in each one?

OpenStudy (anonymous):

yes @mathmate

OpenStudy (mathmate):

 Calculate the volume of water in each glass and the sum should equal 90. Solve for h. Note: if conical glass is full, then the radius is just diameter/2=4.2/2=2.1. Double check my diagram since I don't see the original figure.

OpenStudy (anonymous):

is it 5.62? @mathmate or 4.04?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

What do your calculations give?

OpenStudy (anonymous):

For what?

OpenStudy (anonymous):

i think its 5.62 but idk i need the answer i have like 2 minutes left ;( @mathmate

OpenStudy (mathmate):

 5.6 cm does not look right. The equation to be set up is as follows: \(\pi (4.5/2)^2h + \pi/3(4.2/2)^2(5.6)=90\) Solve for h.